OpenStudy (anonymous):

Use an appropriate formula to write 5*sin 15 degrees * cos 25 degrees- 5*cos 15 degrees * sin 25 degrees as k sin theta or k cos theta, with suitable cvalues for k and theta. Theta must be positive, k can be positive or negative. Do not evaluate. What...

OpenStudy (anonymous):

Use the formula \[ \sin(a-b)= \sin(a) \cos(b) - \cos(a) \sin(b) \]

OpenStudy (anonymous):

So sin(15degrees-25degrees)? but what about the 5 before the first sin and second cosine?

OpenStudy (anonymous):

\[ \sin( 15-25) = \sin(15) \cos(25) - \cos(15) \sin(25)\\ 5\sin( -10) =5 \sin(15) \cos(25) - 5\cos(15) \sin(25)\\ -5\sin(10) =5 \sin(15) \cos(25) - 5\cos(15) \sin(25)\\ k=-5\\ \theta=10 \]

OpenStudy (anonymous):

Multiply the two sides by 5